Evaluate exponential functions Question Let f(x) = (3)+-3. What is f(1)? Provide your answer below: (1)
Answer #1

Given that, f(x) = (3)x-3​​​​​​

So, f(1) = (3)1-3 = (3)-2 = 1/9​​​​​​